Giving a test to a group of students, the grades and gender are summarized below

A B C Total
Male 17 8 5 30
Female 18 19 3 40
Total 35 27 8 70


If one student is chosen at random,

Find the probability that the student was female OR got an "B".
1
Expert's answer
2020-06-02T19:11:39-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 & A & B & C & Total \\ \hline
 Male & 17 & 8 & 5 & 30 \\
 \hdashline
 Female & 18 & 19 & 3 & 40 \\
 \hdashline
 Total & 35 & 27 & 8 & 70
\end{array}"

"P(F\cup B)=P(F)+P(B)-P(F\cap B)="

"={40\over 70}+{27\over 70}-{19\over 70}={24\over 35}"

 The probability that the student was female OR got an "B"is "\dfrac{24}{35}."